# NeuCo-Bench: A Novel Benchmark Framework for Neural Embeddings in Earth Observation

### v1

@@ Abstract
We introduce NeuCo-Bench, a novel benchmark framework for evaluating (lossy) neural compression and representation learning in the context of Earth Observation (EO). Our approach builds on fixed-size embeddings that act as compact, task-agnostic representations applicable to a broad range of downstream tasks. NeuCo-Bench comprises three core components: (i) an evaluation pipeline built around reusable embeddings, (ii) a new challenge mode with a hidden-task leaderboard designed to mitigate pretraining bias, and (iii) a scoring system that balances accuracy and stability. To support reproducibility, we release SSL4EO-S12-downstream, a curated multispectral, multitemporal EO dataset. We present initial results from a public challenge at the 2025 CVPR EARTHVISION workshop and conduct ablations with state-of-the-art foundation models. NeuCo-Bench provides a first step towards community-driven, standardized evaluation of neural embeddings for EO and beyond.
Cartoon summary of the NeuCo-Bench framework. Multi-temporal, multi-modal, and multi-channel inputs $x$ are compressed into fixed-size embeddings $z=E(x)$ by a user-defined encoder $E$. The embeddings are then linearly probed on a diverse set of regression and classification downstream tasks to assess the general-purpose quality of $z$.
@@ 1 Introduction
The rapid growth of visual data, from online media to scientific observation, has made efficient compression a central challenge for storage, transmission, and large-scale analysis (Pouyanfar et al., 2018; Wang et al., 2018a; Gomes et al., 2025). Traditional codecs such as JPEG2000 (Skodras et al., 2001) and more recent learned autoencoders (Ballé et al., 2016) are optimized for pixel-level distortion, largely reflecting human visual perception. However, many machine learning pipelines care less about perceptual fidelity and more about semantic fidelity, retaining the information needed to solve downstream tasks (Huang and Wu, 2024). This gap is particularly critical in domains like Earth Observation (EO), where petabyte-scale datasets of multi-modal satellite imagery must support diverse analytical tasks ranging from environmental monitoring to disaster response (Guo et al., 2017). EO data are characterized by substantial redundancy and noise across multiple spectral bands and temporal sequences, amplifying the need for compression strategies that efficiently capture underlying, task-relevant information (Gomes et al., 2025). This gives rise to the question: How much task‐relevant information can be squeezed into compact data representations?
Recent work has shown that compressed latent representations can preserve rich semantic content, enabling pipelines to operate directly on features without reconstructing the input image (Torfason et al., 2018; Singh et al., 2020). Self-supervised foundation models (FMs) further demonstrate that embeddings can transfer across tasks with minimal fine-tuning. Yet, their dimensionality often rivals or exceeds the size of the original data, reintroducing storage and bandwidth bottlenecks (Gomes and Brunschwiler, 2024; Lu et al., 2024). Despite these advances, there is currently no standardized framework evaluating how effectively compressed representations retain semantic content across multiple downstream tasks. Existing evaluations remain fragmented, often restricted to pixel fidelity, single-task utility, or unconstrained high-dimensional embeddings, making it challenging to compare approaches on a common basis.
@@
  • Section˜5Benchmark Evaluation: We validate the utility of NeuCo-Bench through the 2025 CVPR EARTHVISION data challenge, introducing a novel hidden-task evaluation scheme. We further test embedding quality under diverse compression strategies, including pre-trained neural compressors and FMs.
@@ 2 Related Work
Classical rate-distortion compression. Image and video codecs such as JPEG, JPEG2000, H.264/HEVC (Wallace, 1991; Skodras et al., 2001; Sullivan et al., 2012; Richardson, 2010) exploit handcrafted transforms (Goyal, 2001; Bracewell, 1986; Daubechies, 1992) and entropy coding to reduce statistical redundancy. Their performance is evaluated through the rate–distortion (RD) trade-off between compressed bit rate and reconstruction fidelity (e.g., MSE, PSNR).
@@
Neural image compression. Learned autoencoders replace handcrafted transforms with analysis and synthesis networks jointly optimized for rate and distortion. Differentiable entropy models enable superior RD performance compared to JPEG2000 (Ballé et al., 2016; Theis et al., 2022), with subsequent extensions using hyperpriors (Ballé et al., 2018; Minnen et al., 2018), autoregressive models (Minnen and Singh, 2020), and transformers (Qian et al., 2022). With automated vision pipelines, the concept of compression for machines shifts focus from human‐perceptual fidelity to task‐driven utility. End-to-end approaches jointly optimize compressors with task networks (Chamain et al., 2020, 2021; Le et al., 2021; Codevilla et al., 2021; Wang et al., 2021, 2023a; Fischer et al., 2025). Other methods enforce invariance to task-relevant augmentations through self-supervised objectives (Dubois et al., 2022), or bypass reconstruction by training tasks directly on compressed latents (Torfason et al., 2018; Duan et al., 2023; Singh et al., 2020).
@@
Compression in EO. EO imagery presents unique compression challenges, with multi-spectral bands, temporal sequences, and petabyte-scale archives (Guo et al., 2017; Wilkinson et al., 2024). Traditional pipelines often rely on codecs like JPEG2000 (Yeh et al., 2005). Recent neural approaches extend rate-distortion autoencoders to EO imagery, achieving significant rate-distortion improvements on multispectral data (Alves de Oliveira et al., 2021; Kong et al., 2021; Cao et al., 2022), while temporal compression remains underexplored (Du et al., 2024; Wang et al., 2018b). For a comprehensive review, see (Gomes et al., 2025). Importantly, most works evaluate RD, not task relevance.
@@
EO Foundation Models. Self-supervised learning has enabled large-scale vision foundation models (FMs) pretrained on vast, unlabeled satellite datasets using masked reconstruction, contrastive, or predictive tasks (Wang et al., 2022a; Sun et al., 2022; Wang et al., 2022b; Mai et al., 2022; Wang et al., 2023b; Hong et al., 2023; Jakubik et al., 2023; Liu et al., 2024). These FMs produce versatile high-dimensional embeddings for EO downstream applications, such as flood segmentation, land-use mapping, and environmental monitoring. More recently, multimodal EO foundation models have begun fusing data modalities, such as SAR and optical imagery, to capture diverse geophysical characteristics and improve application performance (Li et al., 2022; Fuller et al., 2023; Xiong et al., 2024; Jakubik et al., 2025; Brown et al., 2025). However, with the exception of (Brown et al., 2025), the resulting latent representations often rival or exceed the original data size, creating data transfer and data processing bottlenecks. Gomes and Brunschwiler (2024) addresses these challenges by integrating neural compression into FM bottlenecks. On image level, Rolf et al. (2021) utilizes fixed, random convolutional kernels to engineer features as basis for linear regression to predict user labels.
@@
EO benchmarks. Current EO domain benchmarks, such as GEO-Bench (Lacoste et al., 2023) and PANGAEA (Marsocci et al., 2025), evaluate FMs by fine-tuning backbones or training complex decoders on intermediate features. These approaches typically require model access and significant computational resources, with limited consideration given to factors such as embedding size and workflow efficiency. In contrast, NeuCo-Bench evaluates fixed-size embeddings through task-agnostic linear probing without any need for access to model backbones. In fact, our approach treats the encoder as a black box that converts any input to a given number of features. NeuCo-Bench provides a lightweight, size-aware evaluation protocol for efficient local testing, and it is structured as a flexible, extendable framework, designed to accommodate future downstream tasks and evaluation methods. Moreover, NeuCo-Bench can be deployed as a novel challenge format that simulates real-world scenarios by requiring participants to submit compressed EO embeddings without prior knowledge of the specific downstream tasks. This setup reflects the demand for broadly generalizable embeddings. As demonstrated in the 2025 CVPR EARTHVISION data challenge, NeuCo-Bench integrates with established platforms such as EvalAI (Yadav et al., ), and is designed to support future competitions on new, unseen tasks.
@@ 3 Benchmarking Framework
Evaluation workflow. Figure˜1 visualizes the pipeline for an Experiment $p$ compressing the samples indexed by $i$ of Benchmark data to create a set of fixed-size embeddings (Embedding set) through an Encoder $E^{(p)}$: These embeddings $X^{(p)}$ are provided to NeuCo-Bench, which performs the evaluation given corresponding Task labels (aka downstream tasks $t=1\dots T$) to return the Benchmark results through a Leaderboard. For each Experiment, the NeuCo-Bench framework performs an Embedding Evaluation given Multimodal, multitemporal datacubes across a set of downstream tasks undisclosed to the developers of a given Compression Model comprising an Experiment $p$. Correspondingly, our framework aggregates scores $s_{t,k}$ per training and test split $k$ to gather statistics for the quality score $Q_{t}$ per downstream task $t$. Consequently, a Scoring algorithm applies a task difficulty-dependent ranking scheme. A standalone Python implementation is available online [^1] https://github.com/embed2scale/NeuCo-Bench. Section˜B.2 provides additional details.
Evaluating embeddings. For NeuCo-Bench, each input sample must be represented as a fixed-size embedding to compress an input data cube, e.g., our EO downstream tasks as detailed in Section˜4, or future extensions, cf. Section˜6. Currently, image-level linear regression and binary (softmax) classification are supported. NeuCo-Bench enforces embeddings of fixed, but configurable, size but otherwise does not constrain how embeddings are generated. Following Fig.˜1, NeuCo-Bench evaluates the compressed embeddings $X^{(p)}$ of an experiment $p$ as follows: For each task $t=1\ldots T$, $K$ linear classifiers (with $N$ tunable parameters $a_{1\dots N}$ plus bias term $a_{0}$) are trained to fit the downstream task labels $y_{\text{ID}}$. Each $k=1\dots K$ denotes a separate, randomly generated split of the downstream task $t$ into a training and testing set. For each tuple $(t,k)$ NeuCo-Bench computes an accuracy measure $s_{t,k}$, utilizing $R^{2}$ (R-squared) for regression tasks and the $F_{1}$ score for (binary) classification. From the set $\left\{s_{t,k}\right\}_{k=1\dots K}$, NeuCo-Bench derives a signal-to-noise-like quality score $Q_{t}^{(p)}$ as the mean performance on task $t$ sensitive to the variability in performance of experiment $p$:
@@
Here, $\left\langle\cdot\right\rangle_{k}$ denotes averaging and $\text{std}_{k}\left(\cdot\right)$ the standard deviation as calculated over the $K$ splits. The parameter $\epsilon>0$ acts as a regulator avoiding high variability in $Q_{t}^{(p)}$ for small $\text{std}_{k}\left(s_{t,k}\right)$. The quality score Eq.˜1 varies in $[0,100]$ for both, classification and regression. Thus, $Q_{t}^{(p)}$ allows for an interpretation of mean accuracy in percent. Compared to using the mean $R^{2}$ over the $K$ splits, $Q_{t}^{(p)}$ penalizes methods with larger variance in the $R^{2}$. Further details on the quality score is provided in Section˜A.1 of the suplpementary material. We note: experiments that perform worse than simply predicting the mean of labels $y_{\text{ID}}$ for regression tasks result in negative $s_{t,k}$ degrading the mean performance. In fact, compression models with negative $Q_{t}^{(p)}$ should be flagged unreliable—they seriously underperform.
Task difficulty-dependent ranking. A novel scoring method is introduced by NeuCo-Bench. It is designed to compare the overall performance of multiple participants over multiple tasks. Based on a rank-then-aggregate approach (Wiesenfarth et al., 2021), NeuCo-Bench dynamically weights the performance across tasks depending on their relative difficulty: Each experiment initially receives a rank $R_{t}^{(p)}$ per task, with the best rank given to the experiment with highest $Q_{t}^{(p)}$. To break ties, all tied experiments are given the lower (better) rank. An experiment’s final rank is calculated from the weighted mean rank across all tasks:
$s^{(p)}=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}\quad\text{with}\quad w_{t}=\text{std}_{p}\left(Q_{t}^{(p)}\right)\Big/\sum_{t=1}^{T}\text{std}_{p}\left(Q_{t}^{(p)}\right)$ (2)
where the tasks are weighted by the standard deviation of the $Q_{t}^{(p)}$ of all experiments on the task. The weighting scales the importance of the tasks such that (a) tasks where all participants perform similarly receive low importance, and (b) tasks where the participants differentiate between each other are weighted highly. NeuCo-Bench also provides the mean $Q$ value $\langle Q_{t}^{(p)}\rangle_{t}$ as an experiment-specific measure of performance. For scenarios with few experiments where the interpretation of a ranking is limited in terms of task difficulty, $\langle Q_{t}^{(p)}\rangle_{t}$ serves as an alternative metric to compare (individual) experiments. Section˜A.1 provides additional analysis of the ranking scheme.
@@ 4 Benchmark Tasks
NeuCo-Bench provides a set of pre-processed, heterogeneous downstream tasks designed for continuous extension in the future. The initial release provides regression labels that are easily turned into binary classification tasks through a threshold.
@@
Dataset
Spatial
Coverage
Temporal
Coverage
Years
of Labels
# Samples
# Tasks
@@
The Earth observation satellite data of the downstream tasks summarize in Table˜1 follow the SSL4EO-S12 data cube structure (Blumenstiel et al., 2025; Wang et al., 2023b). We utilize 13 channels of Sentinel-2 Level-1C Top-of-the-Atmosphere (S2L1C) and 12 channels of multi-spectral Sentinel-2 Level-2A surface reflectance (S2L2A). On top we spatially align 2 channels of radar Sentinel-1 (S1) GRD product polarizations (VV and VH). For a given geolocation we retrieve four timestamps, one per season: winter (Dec–Feb), spring (Mar–May), summer (Jun–Aug), and fall (Sep–Nov). Figure˜1 depicts these four seasonal data cubes, with each containing 27 bands. Google Earth Engine (Gorelick et al., 2017; GEE, ) (GEE) was utilized to download all relevant satellite data. All labels, except for the Clouds use case, have been retrieved from GEE, too. Following (Blumenstiel et al., 2025), the processed data was stored as cloud-ready ZIP-Store of the Zarr file format.
@@
The downstream tasks contain between 1100 and 4691 samples (locations/labels), which are distributed globally (Fig.˜2). The associated satellite data cubes are pre-processed and filtered to ensure UTM-projected patches with a size of 264 × 264 pixels without spatial overlap.
@@
The crops task covers cropland in the US Corn Belt and is provided by the US Department of Agriculture (USDA) (Boryan et al., 2011). Soybean and corn were selected as primary focus classes for this downstream task, with the fraction of corn and soybean within each patch serving as the label. The Crop Data Layer is published annually with a spatial resolution of 30 meters. The labels are available as post-processed data222e.g., the label year 2023 corresponds to crops cultivated during the 2022 growing season.
@@
The landcover tasks leverages aggregated land use data from the European Environment Agency (EEA), which includes various land cover classes such as forests, urban areas, water areas, and agricultural land (European Environment Agency (2018), EEA). These labels represent the dominant land cover within each patch at a spatial resolution of 100 meters within Europe. Based on this data, two downstream tasks are provided for forests and agricultural land 2018 within the challenge.
@@
The biomass tasks uses above-ground biomass estimates derived from LIDAR measurements from the Global Ecosystem Dynamics Investigation (GEDI) instrument. GEDI provides structural information on vegetation height and density, allowing robust models to estimate above-ground biomass in megagrams per hectare (Mg/ha) (Dubayah et al., 2022). Within NeuCo-Bench, the GEDI Level 4A biomass estimates were spatially aggregated to the satellite patches with 264x264 pixels, providing a mean biomass value and its standard deviation as regression targets.
@@
The clouds provides cloud cover fractions based on CloudSen12+ (Aybar et al., 2024) as labels and pre-processed Sentinel-1 and Sentinel-2 data cubes as corresponding observations. Although the SAR data is not affected by clouds, Sentinel-1 is included alongside Sentinel-2 to ensure a consistent data structure for all downstream tasks.
For the heatisland use case, Landsat-8 Land Surface Temperature (LST) provides surface temperature data that are used as labels for urban areas (Observation and Center, 2020). This is particularly relevant in the context of heat events and future urban planning, and contains 1659 samples. The corresponding tasks address the mean surface temperature and its standard deviation per data cube stack. Further information on the data and their downstream tasks during the competition can be found in the Section˜B.1.

### v2

@@ Abstract
We introduce NeuCo-Bench, a novel benchmark framework for evaluating (lossy) neural compression and representation learning in the context of Earth Observation (EO). Our approach builds on fixed-size embeddings that act as compact, task-agnostic representations applicable to a broad range of downstream tasks. NeuCo-Bench comprises three components: (i) an evaluation pipeline built around embeddings, (ii) a challenge mode with a hidden-task leaderboard designed to mitigate pretraining bias, and (iii) a scoring system that balances accuracy and stability. To support reproducibility, we release SSL4EO-S12-downstream, a curated multispectral, multitemporal EO dataset. We present results from a public challenge at the 2025 CVPR EarthVision workshop and conduct ablations with state-of-the-art foundation models. NeuCo-Bench provides a step towards community-driven, standardized evaluation of neural embeddings for EO and beyond.
@@ 1 Introduction
Figure 1: Overview of the benchmark framework. Multi-temporal, multi-modal, and multi-channel inputs $x$ from a spatiotemporal dataset are compressed into fixed-size embeddings $z=E(x)$ by a user-defined encoder $E$. Within the benchmark, these embeddings are linearly probed on a diverse set of regression and classification downstream tasks to assess the general-purpose quality of $z$.
The rapid growth of visual data, from online media to scientific observation, has made efficient compression a central challenge for storage, transmission, and large-scale analysis [50, 73, 27]. Traditional codecs such as JPEG2000 [58] and more recent learned autoencoders [5] are optimized for pixel-level distortion, largely reflecting human visual perception. However, many machine learning pipelines care less about perceptual fidelity and more about semantic fidelity, retaining the information needed to solve downstream tasks [33]. This gap is particularly critical in domains like Earth Observation (EO), where petabyte-scale datasets of multi-modal satellite imagery must support diverse analytical tasks ranging from environmental monitoring to disaster response [30]. EO data are characterized by substantial redundancy and noise across multiple spectral bands and temporal sequences, amplifying the need for compression strategies that efficiently capture underlying, task-relevant information [27]. This gives rise to the question: How much task‐relevant information can be squeezed into compact data representations?
Recent work has shown that compressed latent representations can preserve rich semantic content, enabling pipelines to operate directly on features without reconstructing the input image [64, 56]. Self-supervised foundation models (FMs) further demonstrate that embeddings can transfer across tasks with minimal fine-tuning. Yet, their dimensionality often rivals or exceeds the size of the original data, reintroducing storage and bandwidth bottlenecks [26, 43]. Despite these advances, there is currently no standardized framework evaluating how effectively compressed representations retain semantic content across multiple downstream tasks. Existing evaluations remain fragmented, often restricted to pixel fidelity, single-task utility, or unconstrained high-dimensional embeddings, making it challenging to compare approaches on a common basis.
@@
  • Section˜5Benchmark Evaluation: We validate the utility of NeuCo-Bench through the 2025 CVPR EarthVision data challenge, from here on denoted the EV challenge, introducing a novel hidden-task evaluation scheme. We further test embedding quality under diverse compression strategies, including pre-trained neural compressors and FMs.
@@ 2 Related Work
Image and neural compression. Image codecs such as JPEG, JPEG2000, H.264/HEVC [65, 58, 60, 53] exploit handcrafted transforms [29, 10, 16] and entropy coding to reduce redundancy. Learned autoencoders replace handcrafted transforms with analysis and synthesis networks jointly optimized for rate and distortion. Differentiable entropy models enable superior RD performance compared to JPEG2000 [5, 63], with subsequent extensions using hyperpriors [6, 47], autoregressive models [46], and transformers [51]. Other approaches shift from human‐perceptual fidelity to task‐driven utility by jointly optimizing compressors with task networks end-to-end [13, 14, 39, 15, 67, 68, 24], enforcing invariance to task-relevant augmentations through self-supervised objectives [21], or bypass reconstruction by training tasks directly on compressed latents [64, 19, 56].
@@
Compression in EO. EO imagery presents unique compression challenges, with multi-spectral bands, temporal sequences, and petabyte-scale archives [30, 75]. Traditional pipelines often rely on codecs like JPEG2000 [79]. Recent neural approaches extend rate-distortion autoencoders to EO imagery, achieving significant rate-distortion improvements on multispectral data [3, 37, 12, 18, 69]. Importantly, most works evaluate RD, not task relevance.
@@
EO Foundation Models. Self-supervised learning has enabled vision foundation models (FMs) pretrained on vast, unlabeled satellite datasets using masked reconstruction, contrastive, or predictive tasks [70, 61, 66, 44, 71, 32, 34, 42]. These FMs produce versatile embeddings for EO downstream applications, such as flood segmentation, land-use mapping, and environmental monitoring. More recently, multimodal EO foundation models have begun fusing data modalities, such as SAR and optical imagery, to capture diverse geophysical characteristics and improve application performance [40, 25, 76, 72, 35, 11]. However, the resulting dense representations often rival the original data size, creating transfer and processing bottlenecks. Gomes and Brunschwiler [26] address this by integrating neural compression into FM bottlenecks. At the image level, Rolf et al. [54] utilize fixed, random convolutional kernels to engineer features, while FM embeddings are often obtained from a CLS token or by averaging latent representations.
Implicit neural representations. INRs have emerged as a compelling alternative for compactly encoding Earth observation data. INR-based approaches have been explored for global location embeddings from satellite imagery [36], hyperspectral compression using neural radiance fields [80, 52], and remote sensing image compression via coordinate-based networks [41]. More generally, INRs have shown strong potential for image compression and continuous signal representation [59, 22, 57].
@@
EO benchmarks. Current EO domain benchmarks, such as GEO-Bench [38] and PANGAEA [45], evaluate FMs by fine-tuning backbones or training large decoders on intermediate features. These approaches typically require model access and significant computational resources, with limited consideration given to factors such as embedding size and workflow efficiency. In contrast, NeuCo-Bench evaluates fixed-size embeddings through task-agnostic linear probing without any need for access to model backbones.
Our approach treats the encoder as a black box that converts inputs to a given number of features and provides a lightweight, size-aware evaluation protocol for efficient local testing. It is structured as a flexible, extendable framework, designed to accommodate future downstream tasks (e.g., Copernicus-Bench [72]) and evaluation methods. Moreover, NeuCo-Bench can be deployed as a novel challenge format that simulates real-world scenarios by requiring participants to submit EO embeddings without prior knowledge of the specific downstream tasks. This setup reflects the demand for broadly generalizable embeddings. As demonstrated in the EV challenge, NeuCo-Bench integrates with established platforms such as EvalAI [78], and is designed to support future competitions on new, unseen tasks.
@@ 3 Benchmarking Framework
Evaluation workflow. Figure˜2 visualizes the pipeline for an Experiment $p$ compressing the samples indexed by $i$ of Benchmark data to create a set of fixed-size embeddings (Embedding set) through an Encoder $E^{(p)}$: These embeddings $X^{(p)}$ are provided to NeuCo-Bench, which performs the evaluation given corresponding Task labels (aka downstream tasks $t=1\dots T$) to return the Benchmark results through a Leaderboard. For each Experiment, the benchmark framework performs an Embedding Evaluation given Multimodal, multitemporal datacubes across a set of downstream tasks undisclosed to the developers of a given Compression Model comprising an Experiment $p$. Correspondingly, our framework aggregates scores $s_{t,k}$ per training and test split $k$ to gather statistics for the quality score $Q_{t}$ per downstream task $t$. Consequently, a Scoring algorithm applies a task difficulty-dependent ranking scheme. A standalone Python implementation is available online [^1] https://github.com/embed2scale/NeuCo-Bench. See supplementary material for additional details.
Evaluating embeddings. For NeuCo-Bench, each input sample must be represented as a fixed-size embedding to compress an input data cube, e.g., our EO downstream tasks as detailed in Section˜4, or future extensions, cf. Section˜6. Currently, image-level linear regression and binary (softmax) classification are supported. NeuCo-Bench enforces embeddings of fixed, but configurable, size but otherwise does not constrain how embeddings are generated. The embedding size can differ between tasks. Following Fig.˜2, the benchmark evaluates the compressed embeddings $X^{(p)}$ of an experiment $p$ as follows: For each task $t=1\ldots T$, $K$ linear classifiers (with $N$ tunable parameters $a_{1\dots N}$ plus bias term $a_{0}$) are trained to fit the downstream task labels $y_{\text{ID}}$. Each $k=1\dots K$ denotes a separate, randomly generated split of the downstream task $t$ into a training and testing set. For each tuple $(t,k)$ the benchmark computes an accuracy measure $s_{t,k}$, utilizing $R^{2}$ (R-squared) for regression tasks and the $F_{1}$ score for (binary) classification. From the set $\left\{s_{t,k}\right\}_{k=1\dots K}$, NeuCo-Bench derives a signal-to-noise-like quality score $Q_{t}^{(p)}$ as the mean performance on task $t$ sensitive to the variability in performance of experiment $p$:
@@
Here, $\left\langle\cdot\right\rangle_{k}$ denotes averaging and $\text{std}_{k}\left(\cdot\right)$ the standard deviation as calculated over the $K$ splits. The parameter $\epsilon>0$ acts as a regulator avoiding high variability in $Q_{t}^{(p)}$ for small $\text{std}_{k}\left(s_{t,k}\right)$. The quality score Eq.˜1 varies in $[0,100]$ for both, classification and regression. Thus, $Q_{t}^{(p)}$ allows for an interpretation of mean accuracy in percent. Compared to using the mean $R^{2}$ over the $K$ splits, $Q_{t}^{(p)}$ penalizes methods with larger variance in the $R^{2}$. Further details on the quality score is provided in the supplementary material. We note: experiments that perform worse than simply predicting the mean of labels $y_{\text{ID}}$ for regression tasks result in negative $s_{t,k}$ degrading the mean performance. In fact, compression models with negative $Q_{t}^{(p)}$ should be flagged unreliable—they seriously underperform.
Task difficulty-dependent ranking. A novel scoring method is introduced by NeuCo-Bench. It is designed to compare the overall performance of multiple participants over multiple tasks. Based on a rank-then-aggregate approach [74], the benchmark dynamically weights the performance across tasks depending on their relative difficulty: Each experiment initially receives a rank $R_{t}^{(p)}$ per task, with the best rank given to the experiment with highest $Q_{t}^{(p)}$. To break ties, all tied experiments are given the lower (better) rank. An experiment’s final rank is calculated from the weighted mean rank across all tasks:
$s^{(p)}=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}\quad\text{with}\quad w_{t}=\frac{\text{std}_{p}\left(Q_{t}^{(p)}\right)}{\sum_{t=1}^{T}\text{std}_{p}\left(Q_{t}^{(p)}\right)}$ (2)
where the tasks are weighted by the standard deviation of the $Q_{t}^{(p)}$ of all experiments on the task. The weighting scales the importance of the tasks such that (a) tasks where all participants perform similarly receive low importance, and (b) tasks where the participants differentiate between each other are weighted highly. NeuCo-Bench also provides the mean $Q$ value $\langle Q_{t}^{(p)}\rangle_{t}$ as an experiment-specific measure of performance. For scenarios with few experiments where the interpretation of a ranking is limited in terms of task difficulty, $\langle Q_{t}^{(p)}\rangle_{t}$ serves as an alternative metric to compare (individual) experiments. Based on the mode of operation, $s^{(p)}$ or $Q^{(p)}_{t}$ may be preferred. The former serves competitive challenge settings with novel downstream tasks of unknown levels in difficulty. The latter is favorable for long-term leaderboards where re-ranking of all methods should not depend on the update of a single method. Additional analysis of the ranking scheme is provided in the supplementary material.
@@ 4 Benchmark Tasks
NeuCo-Bench provides a set of pre-processed, heterogeneous EO downstream tasks designed for continuous extension in the future. The initial release provides regression labels that are easily turned into binary classification tasks through a threshold.
@@
Dataset Spatial Coverage Temporal Coverage Years of Labels # Samples # Tasks
@@
The EO satellite data of the downstream tasks, summarized in Table˜1, follow the SSL4EO-S12 data cube structure [8, 71]. We utilize 13 channels of Sentinel-2 Level-1C Top-of-the-Atmosphere (S2L1C) and 12 channels of multi-spectral Sentinel-2 Level-2A surface reflectance (S2L2A). On top we spatially align 2 channels of radar Sentinel-1 (S1) GRD product polarizations (VV and VH). For a given geolocation we retrieve four timestamps, one per season: winter (Dec–Feb), spring (Mar–May), summer (Jun–Aug), and fall (Sep–Nov). Figure˜2 depicts these four seasonal data cubes, with each containing 27 bands. Google Earth Engine [28, 1] (GEE) was utilized to download all relevant satellite data. All labels, except for the Clouds use case, have been retrieved from GEE, too. The processed data was stored as cloud-ready ZIP-Store of the Zarr file format.
@@
The downstream tasks contain between 1100 and 4691 samples (locations/labels), which are distributed globally (Fig.˜3). The associated satellite data cubes are pre-processed and filtered to ensure UTM-projected patches with a size of 264 × 264 pixels without spatial overlap.
@@
The crops task covers cropland in the US Corn Belt and is provided by the US Department of Agriculture (USDA) [9]. Soybean and corn were selected as primary focus classes for this downstream task, with the fraction of corn and soybean within each patch serving as the label. The Crop Data Layer is published annually with a spatial resolution of 30 meters. The labels are available as post-processed data222e.g., the label year 2023 corresponds to crops cultivated during the 2022 growing season.
@@
The landcover tasks leverages aggregated land use data from the European Environment Agency (EEA), which includes various land cover classes such as forests, urban areas, water areas, and agricultural land [23]. These labels represent the dominant land cover within each patch at a spatial resolution of 100 meters within Europe. Based on this data, two downstream tasks are provided for forests and agricultural land 2018 within the challenge.
@@
The biomass tasks uses above-ground biomass estimates derived from LIDAR measurements from the Global Ecosystem Dynamics Investigation (GEDI) instrument, and are sampled based on [55]. GEDI provides structural information on vegetation height and density, allowing robust models to estimate above-ground biomass in megagrams per hectare (Mg/ha) [20]. For the labels, GEDI Level 4A biomass estimates are spatially aggregated to the satellite patches with 264x264 pixels, providing a mean biomass value and its standard deviation as regression targets.
@@
The clouds provides cloud cover fractions based on CloudSen12+ [4] as labels and pre-processed Sentinel-1 and Sentinel-2 data cubes as corresponding observations. Although the SAR data is not affected by clouds, Sentinel-1 is included alongside Sentinel-2 to ensure a consistent data structure for all downstream tasks.
The heatisland task uses the spatio-temporal mean and standard deviations from Landsat-8 [48] Land Surface Temperature (LST) of urban areas as labels. This is particularly relevant in the context of heat events and future urban planning.
Further information on the data and their downstream tasks can be found in the supplementary material.